how many quarts of a 20% saline solution must be added to 8 quarts of a 15% solution to create a solution that is 17% saline
Answer by Paul(988)

20x+15(8-x)=17(8)
20x+120-15x=136
5x=16
x=3.2
8-3.2=4.8. Hence there should be 20% of 3.2 solution added and 4.8 of 15%